OpenStudy (anonymous):

solve 1/x + 3/8 = 1/2. Use complete sentences to explain how to solve this equation

OpenStudy (campbell_st):

subtract 3/8 from both sides which will leave 1/x = 1/8 (since 1/2 = 4/8) x = 8

OpenStudy (anonymous):

ty may i ask u another

OpenStudy (anonymous):

What is the simplified form of 2x+3/x-2 - x+4/x-2?

OpenStudy (campbell_st):

ok.... both fractions have the same denominator so you will have \[\frac{2x + 3 -(x +4)}{x - 2}\] just simplify the numerator

OpenStudy (anonymous):

x-1/x-2

OpenStudy (anonymous):

awesome. i appreciate this. i have tried to figure it out for hours.

OpenStudy (anonymous):

What is the solution to the rational equation 5/x^2+3x+2 - 1/x+2 = 1/3x+3?